The method of tree ring dating gave the following years A.D. for an archaeological excavation site. Assume that the population of x values has an approximately normal distribution.
1201   1229   1264   1257   1268   1316   1275   1317   1275

(a) Use a calculator with mean and standard deviation keys to find the sample mean year x and sample standard deviation s. (Round your answers to the nearest whole number.)
x =   A.D.
s =   yr

(b) Find a 90% confidence interval for the mean of all tree ring dates from this archaeological site. (Round your answers to the nearest whole number.)
lower limit     A.D.
upper limit     A.D.
Answer #1

a)

x = 1267
s = 37

b)

sample mean, xbar = 1267
sample standard deviation, s = 37
sample size, n = 9
degrees of freedom, df = n - 1 = 8

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, tc = t(α/2, df) = 1.86


ME = tc * s/sqrt(n)
ME = 1.86 * 37/sqrt(9)
ME = 22.94

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (1267 - 1.86 * 37/sqrt(9) , 1267 + 1.86 * 37/sqrt(9))
CI = (1244 , 1290)

lower limit = 1244 A.D.
upper limit = 1290 A.D.